# Explain with Visual Keypoints Like a Real Mentor! A Benchmark for Multimodal Solution Explanation

### v3

@@ 4.3 Keypoint-based Explanation Generation
Results.
Table 4 presents the results of the explanation generation task. Although all models perform relatively well in terms of Correctness by aligning their explanations with the problem, many fail to follow the intended reasoning path (Fidelity) or to reference the given visual keypoints (Referencing). OG models generally struggle, whereas the Qwen2.5-VL series performs relatively well, showing improved performance with increasing model size. OM models still fail to generate coherent explanations or follow instructions. In contrast, CG models achieved the highest scores in this explanation task, indicating their strong capabilities in generating well-grounded explanations. As shown in Table 5, human evaluation results show a similar pattern, with strong correlations between human and LLM judgments as indicated by the Spearman coefficients (Zar 2005), which are 0.770 for Correctness, 0.783 for Fidelity, and 0.788 for Referencing, all statistically significant (p$<$0.05). These results suggest that while most open-source models still struggle, CG models and recent OG models like the Qwen2.5-VL series are capable of generating appropriate explanations.
5 Analyses
5.1 Qualitative Analysis
To analyze how the three categories of models differ in their outputs, we examined the results from representative models in each category: Qwen2.5-VL 7B (OG), Math-PUMA (OM), and Gemini 2.0 Flash (CG).
@@ 5.1 Qualitative Analysis
Figure 5 shows VK identification examples from three model categories. The Qwen2.5-VL tries to reason about which option provides the most informative visual keypoints for the task, yet its final selection is wrong. The Math-PUMA shows neither coherent reasoning nor a correct answer. In contrast, the Gemini 2.0 Flash quickly grasps the core of the instruction, analyzes candidate visual keypoints, and selects the option that best supports understanding the problem.
@@
Figure 6 presents explanation examples from three model categories. In this example, Qwen2.5-VL scored 2 for Correctness, 2 for Fidelity, and 3 for Referencing. Although the model is given keypoints, its approach is not aligned with the reference solution. Its explanation references visual cues such as angles, lengths, symbols, and triangles, which indicates partial alignment with the solution image. However, it ultimately misinterprets the problem and produces incorrect reasoning. Math-PUMA received a score of 1 across all metrics, as it only returned the final answer without any supporting explanation. In contrast, Gemini scored 5 in all metrics. It generated an explanation as if the keypoints were embedded in the solution image, following the same reasoning structure as the reference and correctly referring to the same key visual elements throughout the solution.
Model PS Only VKI Only PS $\cap$ VKI
[…]
Math-PUMA 19.7% 15.5% 4.6%
GPT-4o 14.6% 35.9% 11.3%
Gemini 2.0 F 20.5% 34.5% 23.1%
5.2 Problem Solving and VK Identification
Table 6 reports the proportion of cases where models succeed only on Problem-Solving (PS), only on Visual Keypoint Identification (VKI), or on both. Overall, generalist models tend to perform better on VKI than on PS alone, suggesting that most models are more capable of recognizing keypoints assuming the reasoning path is known, rather than solving the problems end-to-end. Notably, Qwen 72B and Gemini achieve over 20% accuracy in cases involving both PS and VKI, demonstrating relatively strong performance. However, since current models still struggle with problem-solving, improving multimodal mathematical reasoning is essential for advancing performance on both PS and VKI.
@@ 5.3 Do MLLMs Attend to Visual Keypoints?
To analyze whether current open-source models can accurately identify visual keypoints in images, we examined the attention maps of OG and OM models on the multimodal solution explanation tasks. Figure 7 illustrates the attention patterns of the OG and OM models. Both models show a tendency to attend to both global and local regions of the input image. However, despite the explicit inclusion of visual keypoints in the options and summary solution of the Visual Keypoint Identification task, neither model effectively focuses on the relevant visual regions. A similar pattern is observed in the Keypoint-based Explanation Generation task. These findings suggest that current open-source models still struggle to visually attend to the keypoints, even when they are essential for understanding or explaining the problem.
6 Related Works
Language Models for Education.
Recent advances in Large Language Models (LLMs) (Brown et al. 2020) have sparked significant interest in educational applications, particularly in personalized problem recommendation (Park et al. 2024), automated tutoring (Chevalier et al. 2024), and the provision of tailored feedback and customized curricula (Hu et al. 2024; Macina et al. 2023; Feng, Wang, and Sun 2023). For effective education, research suggests that combining textual and visual information enhances comprehension and memory more effectively than using text alone (Arcavi 2003; Stylianou 2010; Lee, Park, and Park 2024). As Clark and Paivio (1991) explains, dual representations supply multiple retrieval cues and cultivate richer mental models. Building on this insight, we introduce the multimodal solution explanation task to enable LLMs to offer learners more comprehensive learning opportunities. This task enables the model to pinpoint the visual keypoints essential for students’ comprehension and to generate explanations grounded in those keypoints, thereby delivering more comprehensive educational support and ultimately improving the overall quality of their learning experience.
Multimodal Mathematical Benchmarks.
Current LLMs show strong performance on mathematical problems, making them valuable tools for students (Zhuang et al. 2024; Luo et al. 2025). To evaluate these models, traditional mathematical benchmarks (Cobbe et al. 2021; Hendrycks et al. 2021) have been crucial in assessing reasoning capabilities. With the growing multimodal capabilities of LLMs, benchmarks such as MathVista (Lu et al. 2023), Math-Vision (Wang et al. 2024a), and MathVerse (Zhang et al. 2024) have extended this evaluation to image-based math problems. Recent efforts like OlympiadBench (He et al. 2024) and MM-MATH (Sun et al. 2024) further assess models not just on final answers but also on their reasoning processes. However, most existing benchmarks focus solely on problem-solving, overlooking educational objectives. To address this gap, we introduce ME2, a new benchmark designed to evaluate a model’s multimodal solution explanation capability. Together, ME2 and the multimodal solution explanation task are expected to advance educational research and foster the development of more effective teaching models.
7 Disscusion and Conclusion
We introduce the multimodal solution explanation task, which requires core mathematics-teaching capabilities for generating intuitive and insightful explanations. To benchmark these capabilities, we propose ME2 benchmark, which consists of two subtasks: (1) Visual Keypoint Identification, where models must select visual keypoints that are essential for understanding and explaining the problem, and (2) Keypoint-based Explanation Generation, which requires models to produce explanations grounded in those visual keypoints that support intuitive reasoning.
Our experimental results show that current MLLMs struggle even with the problem-solving aspects of the ME2 tasks. Furthermore, even under the assumption that the model already knows the solution path, similar to a human tutor, we found that most open-source models still fail to accurately identify visual keypoints, with the exception of closed-source models and recent open-source models. Surprisingly, even math-specialized models perform worse in this aspect. The performance gap becomes even more pronounced in the explanation generation task. To perform well on the multimodal solution explanation task, future open-source models may need the following advancements:
  1. 1.
    Enhanced visual grounding mechanisms that link visual elements to multimodal mathematical reasoning.
  2. 2.
    Advanced solution-path tracking capabilities for multimodal explanations.
We believe this work will inspire future research along two key paths: (1) enhancing models’ mathematical reasoning capabilities through improved visual grounding, and (2) leveraging visual grounding in educational contexts to develop AI systems that can serve as real-world mentors, ultimately providing clear, multimodal, and student-friendly explanations. The code and data will be released publicly.
@@ A Benchmark for Multimodal Solution Explanation
Supplementary Material
Appendix A Overview
This supplementary material includes the following sections: (1) Detailed experimental settings used in our benchmark, (2) Details of our benchmark, and (3) Prompts used for model response generation
Appendix B Experiment Details
In this study, experiments were conducted using the LMMs-eval repository (Li et al. 2024). This repository provides a comprehensive framework for evaluating multi-modal models across various tasks.
B.1 Evaluated Models
In all LLM-based evaluations, we used the gpt-4o-2024-08-06 endpoint. When conducting experiments with open-source multimodal models, we leveraged the official implementation codes in conjunction with publicly available weights from the Huggingface Hub  [^1] https://huggingface.co/models. The following model parameters were used for each model:
@@ B.1 Evaluated Models
B.2 Computational Resources
B.3 Human Evaluation Setup
Three evaluators, all holding a bachelor’s or master’s degree in engineering, are assessing AI model outputs for 80 problems. Specifically, they are evaluating the results produced by the Math-PUMA, Qwen2.5-VL, and Gemini 2.0 Flash models, which represent the OM, OG, and CG categories. Each criterion is being rated on a five-point Likert scale.
B.4 Evaluation on automatic metrics
We report additional evaluation results for Keypoint-based Explanation Generation task using automatic metrics, including BLEU, ROUGE, METEOR, and BERTScore (Papineni et al. 2002; Lin 2004; Banerjee and Lavie 2005; Zhang et al. 2019). The detaileed scores can be found in Table A.
Appendix C Benchmark Details
Our ME2 benchmark consists of a total of 17 chapters and 33 sections as shown in Table B.
Appendix D Prompts
This section compiles all the prompts used in our experiments. The prompts shown in Figure A, Figure B, and Figure C are used to generate model outputs for the Solution Recognition toy task, Visual Keypoint Identification, and Keypoint-based Explanation Generation tasks, respectively. For the Keypoint-based Explanation Generation task, model responses are evaluated using the prompts in Figure D.
Appendix E Ethical Considerations
In this paper, we introduce the ME2 benchmark, curated from an in-house mathematics education platform. All materials were reviewed to ensure compliance with copyright regulations. Data annotation was performed by bilingual annotators with undergraduate degrees, ensuring sufficient mathematical problem-solving skills.
@@ Appendix E Ethical Considerations
Figure A: Prompt used for the Visual Keypoint Identification task in the ME2 benchmark. Prompt inputs are boldfaced.
@@
Figure B: Prompt used for the Keypoint-based Explanation Generation task in the ME2 benchmark. It is designed to generate educationally effective explanations for the given math problem. Prompt inputs are boldfaced.
You are a math solver. For the problem below, **your task is ONLY to output the final answer** in one line. **Do NOT provide any explanation, steps, or clarification. Just write the answer.** Problem: As shown in the figure, there are 5 points: A, B, C, D, and E. When selecting two points among them to form straight lines and rays, let the number of straight lines be a and rays be b. Find the value of a + b. Again, only return the final answer. Any additinal text will be considered incorrect.
Figure C: Prompt used for the Solution Recognition toy task in the ME2 benchmark. It is designed to generate an answer to the given math problem. Prompt inputs are boldfaced.
@@ 3. Use of Key Visual Elements
4. Mathematical Accuracy
- Does the model produce the same final answer as the reference?
- If the model’s answer is correct, give a score of **1**.
- If the model’s answer is incorrect, give a score of **0**.
@@ 4. Mathematical Accuracy
Rating: [[x, y, z, w]]
— where ‘x‘ is your score for correctness, ‘y‘ for reference alignment, ‘z‘ for use of visual elements, and ‘w‘ for mathematical accuracy.
Figure D: GPT evaluation prompt used to assess model outputs for the Keypoint-based Explanation Generation task in ME2. Prompt inputs are boldfaced.

### v4

@@ 6.1 Qualitative Analysis
Figure 6: Error analysis for Visual Keypoint Identification: Correct choice rates, (1) cases where the correct element was chosen but referenced incorrectly, (2) choices containing more keypoints than required, and (3) choices containing fewer keypoints than needed.
@@
Figure 5 shows visual keypoint identification examples from three model categories. Qwen2.5-VL attempts to reason about the most informative keypoints but still selects the wrong option. Math-PUMA shows neither coherent reasoning nor a correct answer. In contrast, Gemini 2.0 Flash correctly interprets the instruction, analyzes the candidates, and chooses the most appropriate keypoints. Overall, similar patterns were consistently observed across examples.
To gain a finer-grained understanding of the models, we analyze failure cases on a 10% subset of the dataset, as shown in Figure 6. We categorize errors into three types: (1) electing the correct elements but providing incorrect descriptions, (2) selecting options that contain extra elements, and (3) selecting options with missing required elements. The three models exhibit similar rates for incorrect descriptions and extra elements but differ substantially in missing elements: Math-PUMA shows the highest rate (38%), followed by Qwen2.5-VL (20%). In contrast, Gemini 2.0 Flash achieves the highest accuracy, while Qwen2.5-VL performs moderately, and Math-PUMA remains the least reliable among the three.
@@
Figure 7 presents explanation examples from three model categories. In this example, Qwen2.5-VL scored 2 in Correctness, 2 in Fidelity, and 3 in Referencing. Despite being provided with keypoints, its reasoning diverges from the reference and only partially aligns with the solution image, ultimately leading to an incorrect interpretation. Math-PUMA received a score of 1 across all metrics, as it only returned the final answer without any supporting explanation. In contrast, Gemini scored 5 across all metrics, generating an explanation that mirrored the reference reasoning and consistently referred to the correct key visual elements. Similar to the experimental results in Figure 6, proprietary models generate coherent explanations, whereas open-source models struggle, with math-specialized ones showing almost no reasoning capability.
@@ 6.2 Do MLLMs Attend to Visual Keypoints?
To analyze how current open-source models fail to visually recognize keypoints, we examined the attention maps of both generalist and math-specialized models on the Visual Keypoint Identification task. Figure 8 illustrates the attention patterns, showing that both models tend to attend well to both global and local regions of the input image. However, despite the explicit inclusion of visual keypoints in the options and summary solution, neither model effectively focuses on the relevant visual regions. A similar pattern is observed in the Keypoint-based Explanation Generation task. While current MLLMs are capable of attending to general visual content (zhang2025mllms), our analysis suggests that they still lack robust mathematical visual grounding ability. Improving visual grounding in mathematical contexts will likely be essential for future models to perform well on the multimodal solution explanation task.
7 Conclusion
We introduce the multimodal solution explanation task and the ME2 benchmark, which assess multimodal mathematics‐teaching capabilities through two complementary subtasks: identifying essential visual keypoints and generating explanations grounded in those keypoints. Our experimental results demonstrate that current models struggle with both problem solving and visual keypoint identification, and that this performance gap becomes more pronounced for open-source models in the explanation generation task. Through our analysis, we find that this limitation stems from the lack of math-specific visual grounding and robust visually grounded reasoning. Enhancing these two abilities will be essential for applying MLLMs effectively in educational settings.
8 Ethics Statement
In this paper, we introduce the ME2 benchmark, curated from an in-house mathematics education platform [^1] https://mathpresso.com/en. All materials were reviewed to ensure full compliance with copyright and data usage regulations. Data annotation was conducted by bilingual annotators holding undergraduate degrees in relevant fields, ensuring adequate mathematical and linguistic proficiency.
9 Acknowledgments
This work was partly supported by an Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korean Government (MSIT) (No. RS-2021-II211343, Artificial Intelligence Graduate School Program (Seoul National University), No.RS-2025-02263598, Development of Self-Evolving Embodied AGI Platform Technology through Real-World Experience), the National Research Foundation of Korea(NRF) grant funded by the Korea government(MSIT)(RS-2024-00354218, RS-2024-00353125). We express special thanks to KAIT GPU project. The ICT at Seoul National University provides research facilities for this study.
Explain with Visual Keypoints Like a Real Mentor!
@@ A Benchmark for Multimodal Solution Explanation
Technical Appendix
Appendix A Experiment Details
In this study, experiments were conducted using the LMMs-eval repository  [^2] https://github.com/EvolvingLMMs-Lab/lmms-eval. This repository provides a comprehensive framework for evaluating multi-modal models across various tasks.
A.1 Computational Resources
A.2 Evaluated Models
When conducting experiments with open-source multimodal models, we leveraged the official implementation codes in conjunction with publicly available weights from the Huggingface Hub  [^3] https://huggingface.co/models. The following model parameters were used for each model:
@@ A.2 Evaluated Models
A.3 LLM Evaluation
In all LLM-based evaluations, we used the gpt-4o-2024-08-06 endpoint. For the Keypoint-based Explanation Generation task, we compared the rankings obtained when Math-PUMA and GPT outputs were evaluated separately by Gemini and GPT. The resulting Kendall’s $\tau$ values were 0.90 for Correctness, 0.84 for Fidelity, and 0.92 for Referencing. Although evaluation bias is often a concern when an LLM assesses models from the same family, the high agreement between the GPT-judge and Gemini-judge indicates that no substantial bias is present.
A.4 Human Evaluation
Three evaluators, all holding a bachelor’s or master’s degree in engineering, are assessing AI model outputs for 80 problems. Specifically, they are evaluating the results produced by the Math-PUMA, Qwen2.5-VL, and Gemini 2.0 Flash models, which represent the math-specialized, generalist, and proprietary models categories. Each criterion is being rated on a five-point Likert scale. LLM–human agreement shows strong correlations, as indicated by the Spearman coefficients (0.770 for Correctness, 0.783 for Fidelity, and 0.788 for Referencing; all $p<0.05$). Human–Human agreement, measured by Krippendorff’s $\alpha$, reached 0.696 for Correctness, 0.571 for Fidelity, and 0.612 for Referencing.
A.5 Automatic Metrics Evaluation
We report additional evaluation results for Keypoint-based Explanation Generation task using automatic metrics, including BLEU, ROUGE, METEOR, and BERTScore. The detaileed scores can be found in Table A.
Appendix B Benchmark Details
Our ME2 benchmark consists of a total of 17 chapters and 33 sections as shown in Table B. During dataset validation, we also reviewed the options related to visual keypoint identification and confirmed their consistency and reliability.
Appendix C Ablation on Solution Summary Anchoring
Since a problem may have multiple valid solutions and the corresponding visual keypoints can vary, we provide the model with a solution summary ($T_{s}^{tldr}$) that anchors a single explanatory direction. To examine the effect of this anchoring, we analyze the qualitative differences when the summary is not provided in Figure A. As shown, without anchoring, the model’s explanations often drift toward alternative reasoning paths or focus on irrelevant keypoints.
Figure A: Comparison of explanations generated with and without the solution summary ($T_{s}^{tldr}$). Without anchoring, models often drift to alternative reasoning paths or irrelevant keypoints. (a) Without $T_{s}^{tldr}$; (b) With $T_{s}^{tldr}$.
@@ Appendix C Ablation on Solution Summary Anchoring
Appendix D Prompts
This section compiles all the prompts used in our experiments. The prompts shown in Figure B, Figure C, and Figure D are used to generate model outputs for the Solution Recognition toy task, Visual Keypoint Identification, and Keypoint-based Explanation Generation tasks, respectively. For the Keypoint-based Explanation Generation task, model responses are evaluated using the prompts in Figure E.
@@ Appendix D Prompts
Figure B: Prompt used for the Visual Keypoint Identification task in the ME2 benchmark. Prompt inputs are boldfaced.
@@
Figure C: Prompt used for the Keypoint-based Explanation Generation task in the ME2 benchmark. It is designed to generate educationally effective explanations for the given math problem. Prompt inputs are boldfaced.
You are a math solver. For the problem below, **your task is ONLY to output the final answer** in one line. **Do NOT provide any explanation, steps, or clarification. Just write the answer.** Problem: As shown in the figure, there are 5 points: A, B, C, D, and E. When selecting two points among them to form straight lines and rays, let the number of straight lines be a and rays be b. Find the value of a + b.
Again, only return the final answer. Any additinal text will be considered incorrect.
Figure D: Prompt used for the Solution Recognition toy task in the ME2 benchmark. It is designed to generate an answer to the given math problem. Prompt inputs are boldfaced.
@@ 3. Use of Key Visual Elements
Rating: [[x, y, z]]
— where ‘x‘ is your score for correctness, ‘y‘ for reference alignment, and ‘z‘ for use of visual elements.
Figure E: GPT evaluation prompt used to assess model outputs for the Keypoint-based Explanation Generation task in ME2. Prompt inputs are boldfaced.